name2004-iss4: Logbook Entries

name2004-iss4: SURFACE Messages: 6 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
117 Sat 17-Jul-2004IRT test (6H/6L)ISS4 (Los Mochis)Steve/John
110 Sat 17-Jul-2004IRT Fixed, Maybe! with Questions?ISS4 (Los Mochis)Steve/John
89 Wed 14-Jul-2004IRT Sensor Installed/CheckISS4 (Los Mochis)Steve/John
37 Thu 08-Jul-2004Pyrgeometer removed and taken to Bahia KinoISS4 (Los Mochis)Charlie
36 Thu 08-Jul-2004Pyrgeometer not workingISS4 (Los Mochis)Charlie
11 Tue 29-Jun-2004Surface serial number documentationISS4 (Los Mochis)Mike

117: SURFACE, Site ISS4 (Los Mochis), Sat 17-Jul-2004 14:50:31 GMT, IRT test (6H/6L)

We are testing the effect of swapping 6H and 6L. Tdome and Tcase contacts.
but note there is a mid-level stratus deck obscuring the sun, so it may not
make much difference today.

First switch was made:      14:50
Returned to original state: 15:15

Effect noted:
Before: about 440
During about: 415
After: about 435

We will repeat this test another day at another site...
110: SURFACE, Site ISS4 (Los Mochis), Sat 17-Jul-2004 02:03:59 GMT, IRT Fixed, Maybe! with Questions?
7/16, ~22:00Z Downloaded new NAMEISS4.DLD file: correct coef. for IRT
	The coefficients in the logger for the Tcase 5th order poly value A5
	was off per what Steve Semmer sent us.  Edited campbell program and
	placed the new file where the 'surface start' script will grab it
	for programming into the campbell.
	Old, Saved .csi/.dld files:
	/iss/config/campbell/name2004/NAMEISS4.DLD.SAVE had a5=14.399
	New files:
	/iss/config/campbell/name2004/NAMEISS4.DLD has a5=15.351

Results: Immediately the values 'came-in-line' with roughly what the
	Penasco/Kino readings have been running: ~450w/M^2 in daytime.
	NOTE: this seems really high based on our expectations of what
	it should be, although the consistency with others is tantilizing.

Overall Changes included:
	1) These new coefficients
	2) Earlier swap of Tc/Td that had been wired backwards on 6L/H

	It's possible the wiring at Penasco/Kino/Mochis may now be backwards on
	Td/Tcase.  Evidence:
	1) The readings are now consistent and unexpectedly (to stupid engr.)
	   high at all 3 sites.
	2) I changed the wiring here to match the IRT-Campbell cable as in
	   Kino and changed this 6L/6H
	3) The comments in the .CSI file are somewhat contrary to what's in
	   the code, and what's in the document SteveS sent indicating that
	   the Tdome is on pin-G (sensor) to 6L on wiring panel; and
	   the Tcase is on pin-E (sensor) to 6H on wiring panel.
	   See included lines from the CSI file below.
What to Watch: determine if the signals are too high.
	decide if the code is correct.

Top Comment of CSI file:
;Program:NAMEiss4 [ISS4 mapr]
;Flag Usage:112
;Input Channel Usage:
;1H     WD          2500mV range
;1L     TEMP
;2H     HUM         2500mV range
;2L     PRESS       2500mV range
;3H     Net Rad     250mV range
;3L     Net Rad
;4H     SOL PSP     25mV range
;4L     SOL PSP
;5H     IR  PYG     25mV range
;5L     IR  PYG
;6H     DOME TEMP PYG  250mV fast range
;/* These are comments */

In-Line Code of Program: note this should grab 6L=tdome, 6H=tcase
;This subroutine calculates the temperature for a 10K
;YSI-44031 type thermistor. The data is input in
;locations 11 & 12 and output in the locations 13 (TDOME)
;& 14(TCASE).
; These coefs are for pyg s/n 28809. calc done 3-30-98
; These are to calc dome temperature.
2:  Polynomial (P55)
 1: 1        Reps
 2: 11       X Loc [ TDOME_1   ]
 3: 13       F(X) Loc [ TDOME_C   ]
 4: -42.776  C0
 5: 181.48   C1
 6: -263.90  C2
 7: 227.48   C3
 8: -96.552  C4
 9: 15.622   C5
; These are coefs to calc case temperature.
3:  Polynomial (P55)
 1: 1        Reps
 2: 12       X Loc [ TDOME_2   ]
 3: 14       F(X) Loc [ TCASE_C   ]
 4: -42.819  C0
 5: 181.40   C1
 6: -263.53  C2
 7: 226.86   C3
 8: -96.137  C4
 9: 14.399   C5

89: SURFACE, Site ISS4 (Los Mochis), Wed 14-Jul-2004 20:55:51 GMT, IRT Sensor Installed/Check
IRT Sensor S/N Installed 2:30Z 7/14
Data OK overnight.
Data still toast as expected in day.
Wiring Change - working with wires between 20:00Z until 20:28Z:
   6H wire was loose and easily pulled out...
   6H/6L wired opposite than at Kino/Penasco sites:
     Changed from 6L - E   to 6L - G as others
     Changed from 6H - G   to 6H - E as others
   This may account for some of the 'offset' at night Hal reported....
     ie the Tc/Td were reversed.
However it is Still Bad in Daytime:
   Examining data after 20:28 indicates this was not the problem,
   data is still ~50w/m^2 which is low.
Check Excitation voltage: difficult to check its level, seems going to ~150mV
   but 'digital meter issue.
Output for Tc/Td to gnd ~15mV.  Vdivider 930ohms ok.

Check for problem in dome:
7/14 22:00-22:30 dome covered with box.  Appears to recover although with
	cloud cover that's a factor.  Will try to repeat when there's more
7/15 17:20Z dome covered, partly/mostly sunny and daytime problem evident.
	signal 'recovered' to ~200w/m^2
7/15 19:10Z box removed from dome. When dome was covered, IR quickly dropped from about 200 W/m2 down to about 100 W/m2. Reading then rose back to 200 level over about 10 minutes. Was slowly decreasing to about 175 by 19Z.

37: SURFACE, Site ISS4 (Los Mochis), Thu 08-Jul-2004 18:56:49 GMT, Pyrgeometer removed and taken to Bahia Kino
Charlie and Hal have taken the pyrgeometer (S/N 28809F3) to Bahia Kino to see if it works on that ISS, and where they have a battery to try. 

Cohn and Militzer will bring it back down to Los Mochis during their circuit.
36: SURFACE, Site ISS4 (Los Mochis), Thu 08-Jul-2004 18:53:37 GMT, Pyrgeometer not working
The pyg has not been working since the system was set up. The IR shows a diurnal trace, with values going to zero during the day.
11: SURFACE, Site ISS4 (Los Mochis), Tue 29-Jun-2004 02:41:36 GMT, Surface serial number documentation
Equipment serial numbers
PIR 28809F
PYR 26225F3
NET 92072    top 5.67   bot 10.69
T/RH v4950020
Press 633423
W-HD 05103-LX*243854
Rain 12788-1193